# Singular Value and Frame Decomposition-based Reconstruction for Atmospheric Tomography

### v1

@@ 1 Introduction
Adaptive Optics (AO) is an essential component of ground-based telescopes such as the Extremely Large Telescope (ELT) [ELT_2020] of the European Southern Observatory (ESO), currently under construction in the Atacama desert in Chile. This is because temperature fluctuations in the atmosphere cause turbuluence, which results in wavefront aberrations of the incoming light before it reaches the telescope. If not corrected by an AO system, these aberrations result in blurred images and therefore a severe loss of image quality. The general working principle of an AO system can be summarized as follows [Roddier1999, Roggemann_Welsh_1996, Ellerbroek_Vogel_2009]: First, a wavefront sensor (WFS) is used to measure the wavefront aberration of the incoming light of some reference light source such as a natural guide stars (NGSs). Then, a deformable mirror (DM) located in the light path is adjusted, such that after reflection from this DM, the incoming wavefront is approximately plane and thus aberration free; see Figure 1.1 (left). Since the atmosphere is rapidly changing, this measurement and correction cycle has to be repeated continuously at a frequency of about 500 Hz.
@@ 2 The Atmospheric Tomography Operator
First, let the domain $\Omega_{A}\subset\mathbb{R}^{2}$ represent the telescope aperture, and assume that there are $L$ atmospheric layers, i.e., planes parallel to the aperture $\Omega_{A}$, located at distinct heights $h_{\ell}\in\mathbb{R}_{0}^{+}$ for $\ell=1\,,\dots\,,L$. We assume that the heights are given in ascending order and note that typically $h_{1}=0$. Next, consider $G$ different guide stars with corresponding direction vectors $\alpha_{g}=(\alpha_{g}^{x},\alpha_{g}^{y})\in\mathbb{R}^{2}$ for $g=1\ ,,\dots\,,G$. The vectors $\alpha_{g}$ are such that seen from the center of the telescope aperture, the vectors $(\alpha_{g}^{x},\alpha_{g}^{y},1)\in\mathbb{R}^{3}$ point directly at the corresponding guide stars. Now, assume that the first $G_{\operatorname{NGS}}$ guide stars are NGS, while the remaining $G_{\operatorname{LGS}}$ guide stars are LGS, such that $G=G_{\operatorname{NGS}}+G_{\operatorname{LGS}}$. Then we can define the coefficients
@@
where $h_{\operatorname{LGS}}$ denotes the height of the sodium layer in the atmosphere in which LGS are created (approximately $90$ km). Note that since $h_{\ell}<h_{LGS}$ for all $\ell=1,\ldots,L$, we have $c_{\ell,g}\in(0,1]$. Next, we introduce the domains
@@
We now want to define the atmospheric tomography operator between suitable Lebesgue spaces. For this, we introduce the space $L_{2}(\Omega,\gamma)$, by which we simply mean the classic Lebesgue space $L_{2}(\Omega)$ equipped with the scaled inner product
@@
Let $\phi=(\phi_{\ell})_{\ell=1,\ldots,L}$ denote the atmospheric turbulence and $\varphi=(\varphi_{g})_{g=1,\ldots,G}$ the incoming wavefronts as reconstructed by the WFSs. Then for given turbulence weights $(\gamma_{\ell})_{\ell=1,\dots,L}$ the atmospheric tomography operator can be defined as [Fusco_Conan_Rousset_Mugnier_Michau_2001]
$A:\mathcal{D}(A):=\prod_{\ell=1}^{L}L_{2}(\Omega_{\ell},\gamma_{\ell})\to L_{2}(\Omega_{A})^{G}\,,$ (2.1)
$\qquad\phi=(\phi_{l})_{l=1}^{L}\mapsto\varphi=\left((A_{g}\phi)(r)\right)_{g=1}^{G}:=\left(\sum_{\ell=1}^{L}\phi_{l}(c_{\ell,g}r+\,\alpha_{g}h_{\ell})\right)_{g=1}^{G}\,.$
@@
The atmospheric tomography operator $A$ as defined in (2.1) essentially only sums up the contributions of each turbulence layer in the direction of the guide stars, and the coefficients $c_{\ell,g}$ model the cone effect for LGS as illustrated in Figure 1.3 (right). The weights $\gamma_{\ell}$ are used to place a higher emphasis on layers on which a strong turbulence is expected, and are assumed to satisfy $\sum_{\ell=1}^{L}\gamma_{\ell}=1$. In practice, they are known quantities derived from previously measured turbulence profiles.
@@
$\begin{split}(A^{\ast}\varphi)(r)&=\sum_{g=1}^{G}(A_{g}^{*}\varphi)(r)=\sum_{g=1}^{G}\left(\frac{\gamma_{\ell}}{c_{\ell,g}^{2}}\varphi_{g}\left(\frac{r-\alpha_{g}h_{\ell}}{c_{\ell,g}}\right)I_{\Omega_{A}(\alpha_{g}h_{\ell})}(r)\right)_{l=1}^{L}\,.\end{split}$ (2.2)
It was shown in [Neubauer_Ramlau_2017] that the atmospheric tomography operator $A$ is not compact, and thus an SVD does not necessarily need to exist. Nevertheless, an SVTD can be derived for a “periodic” atmospheric tomography operator on a square domain, which we now introduce in a general Sobolev setting. For this, let $\Omega_{T}=[-T,T]^{2}\subset\mathbb{R}^{2}$ be a square domain with $T$ chosen sufficiently large such that, cf. Figure 2.1,
@@
$c_{l}:=\min_{g=1,\ldots,G}\left\{c_{\ell,g}\right\}\,,\qquad\forall\,\ell=1,\ldots,L\,.$
@@
This amounts to the classic Fourier-series definition of Sobolev spaces [Adams_1970, Natterer_2001], adapted to our specific scaled domains $c_{\ell}\Omega_{T}$ and incorporating the turbulence weights $\gamma_{l}$. With this, the periodic atmospheric tomography operator is defined as
$\begin{split}&\tilde{A}^{(s)}:\prod_{\ell=1}^{L}H^{s}\left(c_{\ell}\Omega_{T},\gamma_{\ell}\right)\rightarrow L_{2}\left(\Omega_{T}\right)^{G}\,,\\ &\qquad\phi=(\phi_{l})_{l=1}^{L}\mapsto\varphi=\left((\tilde{A}^{(s)}_{g}\phi)(r)\right)_{g=1}^{G}:=\left(\sum_{\ell=1}^{L}\phi_{l}(c_{\ell,g}r+\,\alpha_{g}h_{\ell})\right)_{g=1}^{G}\,.\end{split}$
The compactness of this operator now depends on the specific choice of $s$. If $s=0$, then as above it can be shown that $\tilde{A}^{(0)}$ is not compact, and thus an SVD does not necessarily need to exist. However, if $s>0$, then the compactness of the Sobolev embedding operator on the bounded domain $\Omega_{T}$ implies that $\tilde{A}^{(s)}$ is compact, and thus an SVD exists. In the specific case of $s=0$ and $\gamma_{l}=1$ for all $l=1,\ldots,L$, SVTDs for the NGS-only and LGS-only case have been derived in [Neubauer_Ramlau_2017] and [Hubmer_Ramlau_2020], respectively. As we shall see below, these leverage algebraic properties of the scaled exponential functions $w_{jk}$. In particular, it was shown in [Neubauer_Ramlau_2017, Theorem 5.1] that if
$\frac{\alpha_{g}^{x}h_{l}}{T}\in\mathbb{Q}\,,\quad\text{and}\quad\frac{\alpha_{g}^{x}h_{l}}{T}\in\mathbb{Q}\,,\qquad\forall\,g=1,\ldots,G\,,\quad\forall\,l=1,\ldots,L\,,$ (2.6)
@@ 3 Singular Value Type Decompositions
In this section, we derive an SVD for the periodic atmospheric tomography operator $\tilde{A}^{(s)}$ for the case of either NGS-only or LGS-only. In this case, $c_{\ell,g}=c_{\ell}$ and thus
$\begin{split}&\tilde{A}^{(s)}:\mathcal{D}(\tilde{A}^{(s)}):=\prod_{\ell=1}^{L}H^{s}\left(c_{\ell}\Omega_{T},\gamma_{\ell}\right)\rightarrow L_{2}\left(\Omega_{T}\right)^{G}\,,\\ &\qquad\phi=(\phi_{l})_{l=1}^{L}\mapsto\varphi=\left((\tilde{A}^{(s)}_{g}\phi)(r)\right)_{g=1}^{G}=\left(\sum_{\ell=1}^{L}\phi_{l}(c_{\ell}r+\,\alpha_{g}h_{\ell})\right)_{g=1}^{G}\,.\end{split}$ (3.1)
@@
$\phi=\left(\sum_{j,k\in\mathbb{Z}}\phi_{jk,l}\,w^{(s)}_{jk,\ell}\right)_{\ell=1}^{L}\,.$
@@
This proof of this proposition follows the lines of [Hubmer_Ramlau_2020, Proposition 4.1]: From the definition of $\tilde{A}^{(s)}$ and with the coefficient expansion (3.3) it follows that
@@
$\varphi=\sum_{j,k\in\mathbb{Z}}\varphi_{jk}w_{jk}\,,\qquad\text{where}\qquad\varphi_{jk}=\left(\left\langle\,\varphi_{g},w_{jk}\,\right\rangle_{L_{2}(\Omega_{T})^{G}}\right)_{g=1}^{G}\in\mathbb{C}^{G}\,.$
@@
$\begin{split}&\sum_{\ell=1}^{L}\left\langle\,(v^{(s)}_{jk,n})_{\ell}\,w^{(s)}_{jk,\ell},(v^{(s)}_{j^{\prime}k^{\prime},n^{\prime}})_{\ell}\,w^{(s)}_{j^{\prime}k^{\prime},\ell}\,\right\rangle_{H^{s}(c_{\ell}\Omega_{T},\gamma_{\ell})}\\ &\qquad=\sum_{\ell=1}^{L}(v^{(s)}_{jk,n})_{\ell}\,\overline{(v^{(s)}_{j^{\prime}k^{\prime},n^{\prime}})}_{\ell}\left\langle\,w^{(s)}_{jk,\ell},w^{(s)}_{j^{\prime}k^{\prime},\ell}\,\right\rangle_{H^{s}(c_{\ell}\Omega_{T},\gamma_{\ell})}\\ &\qquad=\left(v^{(s)}_{jk,n^{\prime}}\right)^{H}v^{(s)}_{jk,n}\,\delta_{jk,j^{\prime}k^{\prime}}=\delta_{n,n^{\prime}}\,\delta_{jk,j^{\prime}k^{\prime}}\,,\end{split}$
@@
As mentioned before, the above analysis generalizes results of [Neubauer_Ramlau_2017, Hubmer_Ramlau_2020], which were in particular derived for the case $s=0$. The corollary below provides an explicit relation between the general case $s>0$ and this base setting.
@@
Let $s\geq 0$ and let the operators $\tilde{A}^{(s)},\tilde{A}^{(0)}$ be defined as in (2). Furthermore, let $(\sigma^{(s)}_{jk,n},u^{(s)}_{jk,n},v^{(s)}_{jk,n})$ and $(\sigma^{(0)}_{jk,n},u^{(0)}_{jk,n},v^{(0)}_{jk,n})$ denote the singular systems of the matrices $\tilde{A}^{(s)}_{jk}$ and $\tilde{A}^{(0)}_{jk}$ defined in (3.4), respectively. Then it follows that
$(\tilde{A}^{(s)}\phi)(x,y)=\sum_{j,k\in\mathbb{Z}}\left(1+\beta_{\ell,T}|(j,k)|^{2}\right)^{-s}\left(\sum_{n=1}^{r^{(0)}_{jk}}\sigma^{(0)}_{jk,n}\left((v^{(s)}_{jk,n})^{H}\phi_{jk}\right)u^{(0)}_{jk,n}\right)w_{jk}(x,y)\,,$
@@
$({(\tilde{A}^{(s)})}^{\dagger}\varphi)_{\ell}(x,y)=\sum_{j,k\in\mathbb{Z}}\left(1+\beta_{\ell,T}|(j,k)|^{2}\right)^{s}\left(\sum_{n=1}^{r_{jk}(0)}\frac{\left((u^{(0)}_{jk,n})^{H}\varphi_{jk}\right)}{\sigma^{(0)}_{jk,n}}v^{(0)}_{jk,n}\right)_{\ell}w_{jk,\ell}(x,y)\,.$
@@
Note first that due to (2.5) and (3.4) there holds
$\tilde{A}^{(s)}_{jk}=\left(1+\beta_{\ell,T}|(j,k)|^{2}\right)^{-s/2}\tilde{A}^{(0)}_{jk}\,,$
@@
$(\tilde{A}^{(s)}_{jk})^{H}\tilde{A}^{(s)}_{jk}v^{(0)}_{jk,n}$ $=\left(1+\beta_{\ell,T}|(j,k)|^{2}\right)^{-s}(\tilde{A}^{(0)}_{jk})^{H}\tilde{A}^{(0)}_{jk}v^{(0)}_{jk,n}$
$=\left(1+\beta_{\ell,T}|(j,k)|^{2}\right)^{-s}(\sigma^{(0)}_{jk,n})^{2}v^{(0)}_{jk,n}\,,$
and thus $\sigma^{(s)}_{jk,n}=\left(1+\beta_{\ell,T}|(j,k)|^{2}\right)^{-s/2}\sigma^{(0)}_{jk,n}$, $r_{jk}^{(s)}=r_{jk}^{(0)}$ and $v^{(s)}_{jk,n}=v^{(0)}_{jk,n}$. Similarly, we obtain $u^{(s)}_{jk,n}=u^{(0)}_{jk,n}$ which together with (3.5) and (3.8) yields the assertion. ∎
Note that the above corollary not only connects the SVTDs of $\tilde{A}^{(s)}$ for different cases of $s$, but also sheds some additional light on the ill-posedness of the problem. In particular, due to the factors $\left(1+\beta_{\ell,T}|(j,k)|^{2}\right)^{-s}$, which essentially correspond to the singular values of the Sobolev embedding operator (cf. [Hubmer_Sherina_Ramlau_2023]), the inversion of $\tilde{A}^{(s)}$ is ill-posed for each $s>0$. Furthermore, if (2.6) holds, then the degree of ill-posedness is exactly $s/2$, i.e., the same as for the problem of inverting the Sobolev embedding operator. Note that due to a result by Kolmogorov [Kolmogorov], a typical atmospheric turbulence layer is expected to satisfy $s=11/6$. However, if (2.6) is not satisfied, then the degree of ill-posedness may be significantly worse.

### v2

@@ 1 Introduction
Adaptive Optics (AO) is an essential component of ground-based telescopes such as the Extremely Large Telescope (ELT) [ELT_2020] of the European Southern Observatory (ESO), currently under construction in the Atacama desert in Chile. This is because temperature fluctuations in the atmosphere cause turbulence, which results in wavefront aberrations of the incoming light before it reaches the telescope. If not corrected by an AO system, these aberrations result in blurred images and therefore a severe loss of image quality. The general working principle of an AO system can be summarized as follows [Roddier1999, Roggemann_Welsh_1996, Ellerbroek_Vogel_2009]: First, a wavefront sensor (WFS) is used to measure the wavefront aberration of the incoming light of some reference light source such as a natural guide stars (NGS). Then, a deformable mirror (DM) located in the light path is adjusted, such that after reflection from this DM, the incoming wavefront is approximately plane and thus aberration free; see Figure 1.1 (left). Since the atmosphere is rapidly changing, this measurement and correction cycle has to be repeated continuously at a frequency of about 500 Hz.
@@ 2 The Atmospheric Tomography Operator
First, let the domain $\Omega_{A}\subset\mathbb{R}^{2}$ represent the telescope aperture, which typically (but not always) is a circular and symmetric domain centered around the origin. Furthermore, assume that there are $L$ atmospheric layers, i.e., planes parallel to the aperture $\Omega_{A}$, located at distinct heights $h_{\ell}\in\mathbb{R}_{0}^{+}$ for $\ell=1\,,\dots\,,L$. We assume that the heights are given in ascending order and note that typically $h_{1}=0$. Note that all 2-dimensional domains defined here are embedded in $\mathbb{R}^{3}$ by fixing the z-coordinate. Next, consider $G$ different guide stars with corresponding direction vectors $\alpha_{g}=(\alpha_{g}^{x},\alpha_{g}^{y})\in\mathbb{R}^{2}$ for $g=1\ ,,\dots\,,G$. The vectors $\alpha_{g}$ are such that seen from the center of the telescope aperture, the vectors $(\alpha_{g}^{x},\alpha_{g}^{y},1)\in\mathbb{R}^{3}$ point directly at the corresponding guide stars. Now, assume that the first $G_{\operatorname{NGS}}$ guide stars are NGS, while the remaining $G_{\operatorname{LGS}}$ guide stars are LGS, such that $G=G_{\operatorname{NGS}}+G_{\operatorname{LGS}}$. Then we can define the coefficients
@@
where $h_{\operatorname{LGS}}$ denotes the height of the sodium layer in the atmosphere in which LGS are created (approximately $90$ km). The coefficients $c_{\ell,g}$ model the cone effect for LGS as illustrated in Figure 1.3 (right). Note that since $h_{\ell}<h_{LGS}$ for all $\ell=1,\ldots,L$, we have $c_{\ell,g}\in(0,1]$. Next, we introduce the domains
@@
We now want to define the atmospheric tomography operator between suitable Lebesgue spaces. For this, we introduce the space $L_{2}(\Omega,\gamma)$ with $\gamma>0$, by which we simply mean the classic Lebesgue space $L_{2}(\Omega)$ equipped with the scaled inner product
@@
Let $\phi=(\phi_{\ell})_{\ell=1,\ldots,L}$ denote the refractive index variations (a dimensionless quantity), related to temperature fluctuations within the atmosphere, causing atmospheric turbulence, and let $\varphi=(\varphi_{g})_{g=1,\ldots,G}$ denote the incoming wavefronts as reconstructed by the WFSs. Then for given turbulence weights $(\gamma_{\ell})_{\ell=1,\dots,L}$ the atmospheric tomography operator can be defined as [Fusco_Conan_Rousset_Mugnier_Michau_2001]
$A:\mathcal{D}(A):=\prod_{\ell=1}^{L}L_{2}(\Omega_{\ell},\gamma_{\ell})\to L_{2}(\Omega_{A})^{G},\qquad\phi=(\phi_{\ell})_{\ell=1}^{L}\mapsto\varphi=\left((A_{g}\phi)\right)_{g=1}^{G}\,,$ (2.1)
$\qquad(A_{g}\phi)(r):=\sum_{\ell=1}^{L}\phi_{l}(c_{\ell,g}r+\,\alpha_{g}h_{\ell})\,.$
@@
The atmospheric tomography operator $A$ as defined in (2.1) essentially only sums up the contributions of each turbulence layer in the direction of the guide stars. The weights $\gamma_{\ell}$ are used to place a higher emphasis on layers on which a strong turbulence is expected, and are assumed to satisfy $\sum_{\ell=1}^{L}\gamma_{\ell}=1$. In practice, they are known quantities derived from previously measured turbulence profiles. Note that from a mathematical point of view, one may also consider non-constant (i.e. spatially varying) weights $\gamma_{\ell}$, thereby placing a varying emphasis on certain areas within the atmospheric layers. An example are the piecewise constant weights introduced in [Saxenhuber_Ramlau_2016], which are also briefly discussed at the end of Section 4. Since they are not commonly used in practice, we do not consider such spatially varying weights here, but note that our subsequent analysis could be extended to include them as well.
@@
$\begin{split}(A^{\ast}\varphi)(r)&=\sum_{g=1}^{G}(A_{g}^{*}\varphi)(r)=\sum_{g=1}^{G}\left(\frac{\gamma_{\ell}}{c_{\ell,g}^{2}}\varphi_{g}\left(\frac{r-\alpha_{g}h_{\ell}}{c_{\ell,g}}\right)I_{\Omega_{A}(\alpha_{g}h_{\ell})}(r)\right)_{l=1}^{L}\,,\end{split}$ (2.2)
where $I_{\Omega_{A}(\alpha_{g}h_{\ell})}$ denotes the indicator function of the domain $\Omega_{A}(\alpha_{g}h_{\ell})$. It was shown in [Neubauer_Ramlau_2017] that the atmospheric tomography operator $A$ is not compact, and thus an SVD does not necessarily need to exist. Nevertheless, an SVTD can be derived for a “periodic” atmospheric tomography operator on a square domain, which we now introduce in a general Sobolev setting. For this, let $\Omega_{T}=[-T,T]^{2}\subset\mathbb{R}^{2}$ be a square domain with $T$ chosen sufficiently large such that, cf. Figure 2.1,
@@
$c_{\ell}:=\min_{g=1,\ldots,G}\left\{c_{\ell,g}\right\}\,,\qquad\forall\,\ell=1,\ldots,L\,.$
@@
The functions $w_{jk}$ and $w_{jk,\ell}$ defined in (2.5) form orthonormal bases over the spaces $L_{2}(\Omega_{T})$ and $L_{2}(c_{\ell}\Omega_{T},\gamma_{\ell})$, respectively. Note that our definition of $H^{s}\left(c_{\ell}\Omega_{T},\gamma_{\ell}\right)$ amounts to the classic Fourier-series definition of periodic Sobolev spaces (see e.g. [Hubmer_Sherina_Ramlau_2023, Ramlau_Teschke_2004_1]), adapted to our specific scaled domains $c_{\ell}\Omega_{T}$ and incorporating the turbulence weights $\gamma_{\ell}$. For integer order $s\in\mathbb{N}$, the inner product (2.4) is equivalent to the classic Sobolev space inner product [Hubmer_Sherina_Ramlau_2023, Proposition 7.2], while for general $s\in\mathbb{R}$ equivalence holds over the Triebel spaces, which include zero boundary conditions [Natterer_2001]. With this, the periodic atmospheric tomography operator is now defined as
$\begin{split}&\tilde{A}^{(s)}:\prod_{\ell=1}^{L}H^{s}\left(c_{\ell}\Omega_{T},\gamma_{\ell}\right)\rightarrow L_{2}\left(\Omega_{T}\right)^{G}\,,\qquad\phi=(\phi_{\ell})_{\ell=1}^{L}\mapsto\varphi=\left((\tilde{A}^{(s)}_{g}\phi)\right)_{g=1}^{G}\\ &\qquad(\tilde{A}^{(s)}_{g}\phi)(r):=\sum_{\ell=1}^{L}\phi_{\ell}(c_{\ell,g}r+\,\alpha_{g}h_{\ell})\,,\end{split}$
with $0\leq s\in\mathbb{R}$. The compactness of this operator now depends on the specific choice of $s$. If $s=0$, then as above it can be shown that $\tilde{A}^{(0)}$ is not compact, and thus an SVD does not necessarily need to exist. However, if $s>0$, then the compactness of the Sobolev embedding operator on the bounded domain $\Omega_{T}$ implies that $\tilde{A}^{(s)}$ is compact, and thus an SVD exists [Engl_Hanke_Neubauer_1996]. In the specific case of $s=0$ and $\gamma_{\ell}=1$ for all $\ell=1,\ldots,L$, SVTDs for the NGS-only and LGS-only case have been derived in [Neubauer_Ramlau_2017] and [Hubmer_Ramlau_2020], respectively. As we shall see below, these leverage algebraic properties of the scaled exponential functions $w_{jk}$. In particular, it was shown in [Neubauer_Ramlau_2017, Theorem 5.1] that if
$\frac{\alpha_{g}^{x}h_{\ell}}{T}\in\mathbb{Q}\,,\quad\text{and}\quad\frac{\alpha_{g}^{x}h_{\ell}}{T}\in\mathbb{Q}\,,\qquad\forall\,g=1,\ldots,G\,,\quad\forall\,\ell=1,\ldots,L\,,$ (2.6)
@@ 3 Singular Value Type Decompositions
In this section, we derive an SVTD for the periodic atmospheric tomography operator $\tilde{A}^{(s)}$ for the case of either NGS-only or LGS-only. In this case, $c_{\ell,g}=c_{\ell}$ and thus
$\begin{split}&\tilde{A}^{(s)}:\mathcal{D}(\tilde{A}^{(s)}):=\prod_{\ell=1}^{L}H^{s}\left(c_{\ell}\Omega_{T},\gamma_{\ell}\right)\rightarrow L_{2}\left(\Omega_{T}\right)^{G}\,,\quad\phi=(\phi_{\ell})_{\ell=1}^{L}\mapsto\varphi=\left((\tilde{A}^{(s)}_{g}\phi)\right)_{g=1}^{G},\\ &\qquad(\tilde{A}^{(s)}_{g}\phi)(r)=\sum_{\ell=1}^{L}\phi_{\ell}(c_{\ell}r+\,\alpha_{g}h_{\ell})\,.\end{split}$ (3.1)
@@
$\phi=\left(\sum_{j,k\in\mathbb{Z}}\phi_{jk,\ell}\,w^{(s)}_{jk,\ell}\right)_{\ell=1}^{L}\,.$
@@
The proof of this proposition follows the lines of [Hubmer_Ramlau_2020, Proposition 4.1]: From the definition of $\tilde{A}^{(s)}$ and with the coefficient expansion (3.3) it follows that
@@
$\varphi=\sum_{j,k\in\mathbb{Z}}\varphi_{jk}w_{jk}\,,\qquad\text{where}\qquad\varphi_{jk}=\left(\left\langle\,\varphi_{g},w_{jk}\,\right\rangle_{L_{2}(\Omega_{T})}\right)_{g=1}^{G}\in\mathbb{C}^{G}\,.$
@@
$\begin{split}&\sum_{\ell=1}^{L}\left\langle\,(v^{(s)}_{jk,n})_{\ell}\,w^{(s)}_{jk,\ell},(v^{(s)}_{j^{\prime}k^{\prime},n^{\prime}})_{\ell}\,w^{(s)}_{j^{\prime}k^{\prime},\ell}\,\right\rangle_{H^{s}(c_{\ell}\Omega_{T},\gamma_{\ell})}\\ &\qquad=\sum_{\ell=1}^{L}(v^{(s)}_{jk,n})_{\ell}\,\overline{(v^{(s)}_{j^{\prime}k^{\prime},n^{\prime}})}_{\ell}\left\langle\,w^{(s)}_{jk,\ell},w^{(s)}_{j^{\prime}k^{\prime},\ell}\,\right\rangle_{H^{s}(c_{\ell}\Omega_{T},\gamma_{\ell})}\\ &\qquad=\left(v^{(s)}_{j^{\prime}k^{\prime},n^{\prime}}\right)^{H}v^{(s)}_{jk,n}\,\delta_{jk,j^{\prime}k^{\prime}}=\delta_{n,n^{\prime}}\,\delta_{jk,j^{\prime}k^{\prime}}\,,\end{split}$
@@
As mentioned before, the above analysis generalizes results of [Neubauer_Ramlau_2017, Hubmer_Ramlau_2020], which were in particular derived for the case $s=0$. In the case of only NGS, i.e., when $c_{\ell}=1$ for all $\ell=1,\ldots,L$, the corollary below provides an explicit relation between the general case $s>0$ and this base setting.
@@
Let $s\geq 0$, let the operators $\tilde{A}^{(s)},\tilde{A}^{(0)}$ be defined as in (3.1), and let $c_{\ell}=1$ for all $\ell=1,\ldots,L$ (pure NGS case). Furthermore, let $(\sigma^{(s)}_{jk,n},u^{(s)}_{jk,n},v^{(s)}_{jk,n})$ and $(\sigma^{(0)}_{jk,n},u^{(0)}_{jk,n},v^{(0)}_{jk,n})$ denote the singular systems of the matrices $\tilde{A}^{(s)}_{jk}$ and $\tilde{A}^{(0)}_{jk}$ defined in (3.4), respectively. Then $\beta_{T}:=\beta_{\ell,T}$ is independent of $\ell$ and it follows that
$(\tilde{A}^{(s)}\phi)(x,y)=\sum_{j,k\in\mathbb{Z}}\left(1+\beta_{T}|(j,k)|^{2}\right)^{-s}\left(\sum_{n=1}^{r^{(0)}_{jk}}\sigma^{(0)}_{jk,n}\left((v^{(s)}_{jk,n})^{H}\phi_{jk}\right)u^{(0)}_{jk,n}\right)w_{jk}(x,y)\,,$
@@
$({(\tilde{A}^{(s)})}^{\dagger}\varphi)_{\ell}(x,y)=\sum_{j,k\in\mathbb{Z}}\left(1+\beta_{T}|(j,k)|^{2}\right)^{s}\left(\sum_{n=1}^{r^{(0)}_{jk}}\frac{\left((u^{(0)}_{jk,n})^{H}\varphi_{jk}\right)}{\sigma^{(0)}_{jk,n}}v^{(0)}_{jk,n}\right)_{\ell}w_{jk,\ell}(x,y)\,.$
@@
Note that due to (3.2) and (3.4) there holds
$\tilde{A}^{(s)}_{jk}=\left(1+\beta_{T}|(j,k)|^{2}\right)^{-s/2}\tilde{A}^{(0)}_{jk}\,,$
@@
$(\tilde{A}^{(s)}_{jk})^{H}\tilde{A}^{(s)}_{jk}v^{(0)}_{jk,n}$ $=\left(1+\beta_{T}|(j,k)|^{2}\right)^{-s}(\tilde{A}^{(0)}_{jk})^{H}\tilde{A}^{(0)}_{jk}v^{(0)}_{jk,n}$
$=\left(1+\beta_{T}|(j,k)|^{2}\right)^{-s}(\sigma^{(0)}_{jk,n})^{2}v^{(0)}_{jk,n}\,,$
and thus $\sigma^{(s)}_{jk,n}=\left(1+\beta_{T}|(j,k)|^{2}\right)^{-s/2}\sigma^{(0)}_{jk,n}$, $r_{jk}^{(s)}=r_{jk}^{(0)}$ and $v^{(s)}_{jk,n}=v^{(0)}_{jk,n}$. Similarly, we obtain $u^{(s)}_{jk,n}=u^{(0)}_{jk,n}$ which together with (3.5) and (3.8) yields the assertion. ∎
Note that the above corollary not only connects the SVTDs of $\tilde{A}^{(s)}$ in the pure NGS-case for different values of $s$, but also sheds some additional light on the ill-posedness of the problem. In particular, due to the factors $\left(1+\beta_{T}|(j,k)|^{2}\right)^{-s}$, which essentially correspond to the singular values of the Sobolev embedding operator (cf. [Hubmer_Sherina_Ramlau_2023]), the inversion of $\tilde{A}^{(s)}$ is ill-posed for each $s>0$. Furthermore, if (2.6) holds, then the degree of ill-posedness is exactly $s/2$, i.e., the same as for the problem of inverting the Sobolev embedding operator. However, if (2.6) is not satisfied, then the degree of ill-posedness may be significantly worse. Note that due to a result by Kolmogorov [Kolmogorov], a typical atmospheric turbulence layer is expected to satisfy $s=11/6$. Note that this assumption is commonly used both for the simulation of atmospheric turbulence profiles [Octopus, octopus06] as well as in modern reconstruction algorithms for atmospheric tomography; see e.g. [Auzinger_2017, Eslitzbichler2013, Yudytskiy_Helin_Ramlau_2014].